OpenStudy (anonymous):

Braden improves his score with each practice quiz he takes. The number of incorrect answers can be represented by the expression 30 - 6n where n represents the number of practice quizzes he has taken. How many answers will he get incorrect after taking 5 practice quizzes? 0

OpenStudy (anonymous):

n= no of practice quizzes he took It's given he took 4 quizzes so n=4 30−6×4 =30−24=36 Answer: 36 incorrect